# Towards White Box Deep Learning

### v2

@@ 4.2 Quantitative results
We don’t bother measuring the clean test metrics as it is a flawed approach that tells little about the true generalisation capability of the model. Every metric reported here is computed under strong adversarial regime.
The model achieves ~92% accuracy under AutoAttack [4] (with default parameters: norm=’inf’, eps=0.3, step_size=0.1). The metrics are almost identical for the much faster 40-step PGD Attack. Samples are usually misclassified with low confidence and therefore we compute a reliability curve - a precision-recall curve where the positive class is defined as the set of correctly classified samples (this means that the precision for 100% recall is exactly equal to the accuracy). It turns out that for 80% recall model achieves human-level 98% adversarial precision (Figure 1). This means that in practice human intervention would be required only for the easily detected 20% of adversarial samples to achieve human level reliability of the entire system (and most real-life samples are not adversarial).
@@ 4.3.1 Two Step Layer
Figure A shows the initial and learned shapes of the Two Step function. The function smoothly thresholds input at $0.365$ and then at $0.556$ which means that the interval $[0,0.365]$ is treated as OFF, $[0.365,0.556]$ as MEH and $[0.556,1]$ as ON. Bear in mind that this is not the final decision regarding the pixel state - it’s just the input to the next layer.
Figure 2: Two Step layer - initial and learned.
4.3.2 Convolutional Layer
Quick inspection of the Figure 3 shows that the Convolutional Layer has indeed learned meaningful filters that check if the central pixel lays inside an adequately structured region - intuitively the matching value corresponds to the degree the model considers the pixel as being ON.
Figure 3: Learned convolutional kernels in 32 predefined poses.
@@ 4.3.3 Affine Layer
Figure 4 shows the learned affine base features. Figure B in the Appendix B shows learned features in 32 learned feature-specific poses. It’s hard to imagine substantially more meaningful features.
Figure 4: Learned affine base features.
@@ 4.3.4 Logical Layer
4.4 Decision boundary
Figure 5 shows the minimum perturbations required to change model’s predictions under Boundary Attack [2] for sample test images. The added perturbations are easily spotted and understood by humans. Note that by definition model is extremely unconfident on those samples and in practice the perturbations required to fool the entire system would have to be much larger.
Figure 5: Boundary Attack: minimal perturbations that flip model prediction (model confidence at around 50%).
Figure 6: Visualization of test data as "seen" by the first 3 consecutive layers.
Figure 6 shows how the data is "seen" by the first 3 layers. The consecutive rows and labels denote respectively:
  1. 1.
    adversarial input and predicted class probabilities
  2. 2.
    output of Two Step Layer
  3. 3.
    output of Convolutional Layer
  4. 4.
    top pose of affine features predicting label "3" (with the match value)141414these rows are actually overlaid with the middle row to make it easier to see what has been matched.
  5. 5.
    top pose of affine features predicting label "5" (with the match value)14
Note that in Figure 6 there are 3 misclassified examples and for every one of them it’s easy to understand why the model made the mistake.
@@ 4.5 Ablation study
Most of the heavy lifting regarding pixel-intensity alignment is done by the Two Step Layer which lifts adversarial robustness from 0% to ~86%. The Convolutional Layer stabilizes the training and boosts adversarial robustness by further ~6%pt.
@@ 5 Discussion
Affine Layer is the core component of the model and the main reason for its interpretability. However if applied directly to the input it is easily fooled by attacks. It’s white box nature makes it easy to understand where this vulnerability comes from and to fix it: quick inspection of the mismatched poses as in the Figure 6 reveals that it is predominantly due to the large attack budget of $0.3$ per pixel (which is $30\%$ of the entire allowed pixel intensity range) combined with the large linear regions of the function computed by this layer (due to small amount of poses per feature). Human eye appears to ignore such considerate changes in pixel intensities which means that the naive linear representation of pixel brightness is not semantically adequate. This is exactly how (and why) the first two layers of the model were designed - to fix the inherent vulnerability of otherwise perfectly interpretable affine features. These layers are themselves compatible with the framework of semantic features which suggests a large untapped potential of the notion.
The Logical Layer has been developed as an attempt to formalize the idea of an abstract hidden layer sparse activation structure and in that regard it is the most inspiring; however it might turn out to be too generic (on MNIST the properly-initialized linear probe seems to be sufficient).
@@ 6 Further research
Future work would have to involve more complex datasets and more nuanced semantic features. For example notice the analogy between affine semantic features and the SIFT [10] algorithm - the learnable pose sampling may be interpreted as finding Keypoints and their Descriptors, while previous layers (Two Step and Convolutional) are essentially performing input normalization. This might be a helpful analogy for designing more advanced semantic features that sample across varying scales, sharpness or illumination levels.
There might be a concern that adding more layers to the network would reduce its interpretability. However the role of additional layers would be to widen the receptive field of the network. The Affine Layer defined earlier matches parts-at-locations of center-aligned objects. Adding additional layer would involve a sliding window approach (or Keypoint Detection similar to that in SIFT) to match center-aligned objects inside the window and integrate this with the approximate global location of the window; then the Global Affine Layer matches objects-at-locations which form a scene (a higher-level object). This would in fact be similar to applying Affine Layer to the "zoomed-out" version of the input.
Other ideas for further work include:
  • use semantic features in self-supervised setting to obtain interpretable dimensionality reduction;
  • study non-affine perturbations of features, e.g. projective perturbations;
  • find adequate semantic topology for color space;
  • define semantic features for sound (inspiration for semantic invariants could be drawn from music theory);
  • define semantic features for text, e.g. capture semantic invariance of permutations of words;
  • study more complex logical semantic features in more elaborate scenarios;
  • figure out a way to make logical semantic features differentiable, i.e. make $\mathbf{P}$ learnable;
@@ 7 Acknowledgments
This is work was done as an independent self-funded research.
@@
I thank Piotr Wodziński, Adam Szummer, Tomasz Steifer, Dmytro Mishkin, Andrzej Pacuk and members of the reddit community for the great early feedback on the first version of this paper.
@@ Appendix C Preliminary results for the entire MNIST
Figure 7: Affine base features learned on full MNIST.
Figure 8: Masked non-negative weights of linear probe layer - per output class.
Figure 9: Sum of affine features weighted by positive weights of linear probe - per output class.

### v4

@@ 6.2 Quantitative results
The model achieves ~92% adversarial test accuracy under AutoAttack [9] (with the default parameters: norm=’inf’, eps=0.3, step_size=0.1). The results are almost identical for the much faster 40-step PGD Attack with the same parameters. Adversarial samples are usually misclassified with low confidence and therefore the reliability curve is computed for the PGD-attacked test images, i.e. a precision-recall curve where the positive class is defined as the set of correctly classified samples (this means that the precision for 100% recall is exactly equal to the accuracy). It turns out that for 80% adversarial recall model achieves human-level 98% adversarial precision (Figure 4). This means that in practice human intervention would be required only for the easily detected ~20% of adversarial samples to achieve human-level reliability of the entire system (and most real-life samples are not adversarial).
The clean test accuracy is ~99.5%.
@@ 6.3.1 Two Step Layer
Figure 5 shows the initial and learned shapes of the Two Step function. The function smoothly thresholds input at $0.365$ and then at $0.556$ which means that the interval $[0,0.365]$ is treated as OFF, $[0.365,0.556]$ as MEH and $[0.556,1]$ as ON. Bear in mind that this is not the final decision regarding the pixel state - it’s just the input to the next layer.
Figure 5: Two Step layer - initial and learned. The visible smooth thresholding essentially means that the layer has learned the permissible perturbations (i.e. intervals) of real semantic features corresponding to the state of being OFF (number -1) and ON (number 1).
6.3.2 Convolutional Semantic Layer
Quick inspection of the Figure 6 shows that the Convolutional Semantic Layer has indeed learned meaningful filters that check if the central pixel lays inside an adequately structured region.
Figure 6: Learned convolutional kernels in 32 predefined poses. Convolutional Semantic Layer acts pixel-wise and matches (convolves) these rotated kernels against the $5\times 5$ neighbourhood of a given pixel, returning the maximum matching value. Intuitively, this corresponds to the degree of pixel laying on a "bright line".
@@ 6.3.3 Affine Layer
Figure 7 shows the learned affine base features. Appendix B shows the 8 learned affine features in all of the 32 learned feature-specific poses. It’s hard to imagine substantially more meaningful features.
Figure 7: Learned affine base features. The weight-sharing strategy imposed by the affine locality function results in easily interpretable features.
@@ 6.3.4 Logical Layer
Figure 8: Minimally perturbated images under Boundary Attack [6] to change model’s predictions - for sample test images. The perturbations are often meaningful to humans and introduce semantic ambiguity. Note that by definition of the Boundary Attack the model is extremely unconfident on those samples (model confidence at ~50%) and in practice the perturbations required to fool the entire system would have to be much larger.
Figure 9: Visualization of test data as "seen" by the first 3 layers. The consecutive rows and labels denote respectively: adversarial input and predicted class probabilities; output of Two Step Layer; output of Convolutional Semantic Layer; top pose of affine features predicting label "3" (with the match value above); same for label "5". The last two rows are overlaid with the middle row to make it easier to see what has been matched. Note that there are 3 misclassified examples (first, 8th and last) and for every one of them it’s easy to understand why the model made the mistake: for the first one the model recognizes the middle part of "3" as the top part of "5"; the "5" in the 8th image has specific curve at the top which resembles the top part of "3"; the "3" in the last example has very thin top part which gets removed by the first 2 layers thus creating an ambiguity.
@@ 6.4 Ablation study
Most of the heavy lifting regarding pixel-intensity alignment is done by the Two Step Layer which lifts adversarial robustness from 0% to ~86%. The Convolutional Semantic Layer stabilizes the training and boosts adversarial robustness by further ~6%pt.
@@ 7 Discussion
7.1 Defending against adversaries
Affine Layer is the core component of the model and the main reason for its interpretability. However if applied directly to the input it is easily fooled by attacks. Its white box nature makes it easy to understand where this vulnerability comes from and to fix it: quick inspection of the mismatched poses as in the Figure 9 reveals that it is predominantly due to the large attack budget of $0.3$ per pixel (which is $30\%$ of the entire allowed pixel intensity range) combined with the large linear regions of the function computed by this layer (due to small amount of poses per feature). Human eye appears to ignore such considerate changes in pixel intensities which means that the naive linear representation of pixel brightness is not semantically adequate. This is exactly how (and why) the first two layers of the model were designed - to fix the inherent vulnerability of otherwise perfectly interpretable affine features. These layers are themselves compatible with the framework of semantic features which is an early indication of a large potential of the notion.
7.2 Limitations
Semantic features require locality engineering as they rely on the adequate domain-specific definition of the locality function. Therefore they won’t be applicable in contexts where such locality is hard to define. On the other hand it opens up interesting areas for research.
The Logical Layer has been developed as an attempt to formalize the idea of how the generic hidden layer activation pattern should look like; however it might turn out to be prone to the "no free lunch" theorem (e.g. on MNIST the properly-initialized linear probe seems to be sufficient).
Despite the generality of the notion and the ideas given in Section 8 there is a risk that the white box qualities of semantic features will be difficult to achieve for harder datasets. The preliminary results on CIFAR10 seem promising but also reveal the need for careful engineering of the locality function, see Section 8.
@@ 8 Further research
Future work would have to involve more complex datasets and more nuanced semantic features. For example notice the analogy between affine semantic features and the SIFT [21] algorithm - the learnable pose sampling may be interpreted as finding Keypoints and their Descriptors, while previous layers (Two Step and Convolutional Semantic) are essentially performing input normalization. This might be a helpful analogy for designing more advanced semantic features that sample across varying scales, sharpness or illumination levels.
There might be a concern that adding more layers to the network would reduce its interpretability. However the role of additional layers would be to widen the receptive field of the network. The Affine Layer defined earlier matches shapes-at-locations of center-aligned objects. Adding additional layer would involve a sliding window approach (or Keypoint Detection similar to that in SIFT) to match center-aligned objects inside the window and integrate this with the approximate global location of the window; then the Global Affine Layer matches objects-at-locations which form a scene (a higher-level object). This would in fact be similar to applying Affine Layer to the "zoomed-out" version of the input.
The ablations in Section 6.4 suggest that adversarial vulnerability of DNNs might be mostly due to the fragility of the initial layers. Therefore another direction forward would be to re-examine the known approaches to XAI and try to "bug fix" their encoder networks in similar fashion.
Other ideas for further work on semantic features include:
  • use semantic features in self-supervised setting to obtain interpretable dimensionality reduction
  • study non-affine perturbations of features, e.g. projective perturbations
  • find adequate semantic topology for color space
  • semantic features for sound (inspiration for semantic invariants could be drawn from the music theory)
  • semantic features for text, e.g. semantic invariants of permutations of words/sentences capturing thoughts
  • study more complex logical semantic features in more elaborate scenarios
  • figure out a way to make logical semantic features differentiable, i.e. make $\mathbf{P}$ learnable
@@ 9 Acknowledgments
This is work has been done as an independent self-funded research.
@@
I thank Piotr Wodziński, Adam Szummer, Tomasz Steifer, Dmytro Mishkin, Andrzej Pacuk, Bartosz Zieliński, Hubert Baniecki and members of the reddit community for the great feedback on the initial versions of this paper.
@@ Appendix C Preliminary results for the entire MNIST
Figure 10: Affine base features learned on full MNIST. Due to the learnable feature-specific $(x,y)$-location these can be considered shapes-at-locations.
Figure 11: Masked positive weights of linear probe - per output class (here the dark-blue squares are zeros). There are only few dark red neurons per class which means that the digit representation can be considered sparse.
Figure 12: Sum of base affine features weighted by positive weights of linear probe - per output class. The similarity to actual digits means that the model decomposes the digits into human-aligned shapes-at-locations.